# bibliography.bib

### v2

@@
Abstract.
In Outline of a Theory of Truth, Kripke introduces some of the central concepts of the logical study of truth and paradox. He informally defines some of these—such as groundedness and paradoxicality—using modal locutions. We introduce a modal language for regimenting these informal definitions. Though groundedness and paradoxicality are expressible in the modal language, we prove that intrinsicality—which Kripke emphasizes but does not define modally—is not. We characterize the modally definable relations and completely axiomatize the modal semantics.
Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, Andrew Lee, and Matt Mandelkern for comments on drafts.
1. Introduction
@@
Yet Kripke does not describe the intrinsic sentences using modal locutions. This is a striking contrast. Is it a mere coincidence that modal locutions appear in the description of the grounded and paradoxical sentences but not of the intrinsic sentences? Or does it reflect some inherent discrepancy between the objects being defined?
The standard way of studying these objects obviates this question. In particular, Kripke replaces the informal modal glosses on paradoxicality and groundedness with official extensional mathematical definitions. Nevertheless, the modal glosses are nice examples of modality occurring naturally in a logico-mathematical context, so it is worth regimenting them in a way that retains their modal flavor. For the purposes of taking these modal definitions “at face value,” we introduce a modal language and an attendant interpretation $\mathbf{FIX}$. Our modal language is propositional, but it contains structured atomic formulas (like in first-order logic) as opposed to propositional variables. In particular, there are atomic formulas of the form:
@@
We also use $N(x)$ is an abbreviation for $\neg T(x)\wedge\neg F(x)$. Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting.
We provide a complete and decidable axiomatization of the modal semantics. We introduce the following axioms:
@@
and define $\mathbf{S5[Con,Ground,Min_{n}]}$ as the normal modal logic that results from adjoining these axioms to $\mathbf{S5}$. Our main completeness theorem is the following:
@@
This completeness theorem falls out of an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$. From this characterization, we infer that the set of intrinsic sentences is not modally definable. Thus, it is not a mere coincidence that Kripke does not lapse into modal language when defining intrinsicality. Indeed, we are able to show that the intrinsic sentences are indiscernible within the modal language from another set, namely, the set of inevitable sentences, which we define later in the paper. The characterization of the modally definable sets builds on a classic normal form theorem for $\mathbf{S5}$ from [carnap1946modalities]. According to Carnap’s theorem, every formula is $\mathbf{S5}$-equivalent to a disjunction of formulas in a very special form. This normal form result and its corollaries adumbrated here resemble classic results in model theory (particular, results in o-minimality and quantifier elimination). We conclude with a discussion of this analogy.
2. Kripke’s Theory of Truth
@@
2.1. Strong Kleene logic
@@ 2.2. Classification of sentences
  • $\lambda$ has the same semantic value as $\neg\mathsf{True}(\ulcorner\lambda\urcorner)$ in every fixed point.
  • $\tau$ has the same semantic value as $\mathsf{True}(\ulcorner\tau\urcorner)$ in every fixed point.
  • $\gamma$ has the same semantic value as $\mathsf{True}(\ulcorner\gamma\urcorner)\vee\neg\mathsf{True}(\ulcorner\gamma\urcorner)$ in every fixed point.
@@ 3. Modal Systems
4. Single Variable Definability
@@ 4.5. Indiscernibility
    $\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
@@
    $\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
@@
    $\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
@@
    $\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\neg\lozenge N(x)$
@@
    $\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\neg\lozenge N(x)$
@@
    $\neg\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
Note that the set of inevitable sentences is defined by the disjunction $(ii)\vee(iii)\vee(iv)\vee(v)$.
Suppose, for the sake of contradiction, that (2) fails. Then there is some $\star$ such that $y^{\star}=\psi$ is inevitable and that $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(y)$. So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., (ii)–(iv). Let’s break into cases.
Case 1: $\varphi$ is missing disjunct (ii). But $\gamma$ is intrinsic and $ii(\gamma)$. So if $x^{\star}=\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 2: $\varphi$ is missing disjunct (iii). But $\neg\gamma$ is intrinsic and $iii(\neg\gamma)$. So if $x^{\star}=\neg\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 3: $\varphi$ is missing disjunct (iv). But $0=0$ is intrinsic and $iv(0=0)$. So if $x^{\star}=0=0$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 4: $\varphi$ is missing disjunct (v). But $0=1$ is intrinsic and $v(0=1)$. So if $x^{\star}=0=1$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$. ∎
@@ 5.1. Counting Formulas
5.1.1. The Bare Modal System
@@
5.1.2. With the Consistency Axiom
@@ 5.2. Matrices
Table 1. All extensional 2-isolators.
@@
Table 2. Standard numbering of the $3\times 3$ matrix.
@@
Table 3. A proper subset of the extensional 2-isolators.
@@
Table 4. A proper subset of the extensional 2-isolators.
@@
Table 5. A proper subset of the extensional 2-isolators.
@@
Table 6. A proper subset of the extensional 2-isolators.
@@ 5.5. Completeness
5.6. Counting Definable Relations
@@
6. Conclusions
The results in this paper concern extensions of the modal logic $\mathbf{S5}$. Many of the interesting phenomena that arise in modal logic do not arise in $\mathbf{S5}$. $\mathbf{S5}$ has Kripke semantics, but the accessibility relation is not so interesting: Often we think of the worlds as just forming an unstructured set (as we do in this paper). Moreover, as [carnap1946modalities] observed, neighboring modalities collapse into one modal, so there is no non-trivial iteration of modality in $\mathbf{S5}$. So a pessimist might conclude that $\mathbf{S5}$ is somewhat boring. The results presented in this paper suggest a different perspective, namely, that by precluding some interesting phenomena, $\mathbf{S5}$ makes room for others.
In support of this point, let me sketch an analogy with first-order model theory. Carnap’s normal form theorem resembles some classic results of first-order model theory. Indeed, normal form theorems and attendant characterizations of definable sets are standard goals in model theory when they are possible. Recall that model theorists often divide first-order theories into the tame and the wild. The following contrasting theorems give some sense of the divide between the tame and the wild:
Theorem 6.1 ([godel1931formal]).
The first-order theory of $\mathbb{Z}$ as a ring is not decidable.
Theorem 6.2 ([robinson1949definability]).
The first-order theory of $\mathbb{Q}$ as a field is not decidable.
Theorem 6.3 (Tarski 1930).
The first-order theory of $\mathbb{R}$ as a field is decidable. [^2] See [tarski1998decision].
The decidable/undecidable dichotomy is not exactly the same as the model-theoretic tame/wild dichotomy, but in these particular instances they coincide. In particular, Tarski’s theorem follows from o-minimality. A structure is o-minimal if every definable set is a finite union of intervals333‘Interval’ is understood in a broad sense to include rays and points. and a theory is o-minimal if all its structures are. Of course, there is much that cannot be expressed over o-minimal structures. The first-order language of rings, for instance, cannot define the set of integers or the sine function over the reals. However, this permits logicians to give sharp, detailed characterizations of definable sets.
@@ 6. Conclusions
Sometimes wild theories like arithmetic have fragments that are both tame and rich. One method for isolating tame fragments of wild theories is pruning the signature (i.e., prohibiting some symbols from appearing in the formulas). A famous instance of this is Presburger’s theorem (contrast with Gödel’s theorem):
Theorem 6.4 ([presburger1929uber]).
The theory of $\mathbb{Z}$ as an additive group is decidable.
Presburger’s theory is tame according to various model-theoretic criteria, e.g., it eliminates quantifiers.
Another method for isolating tame fragments of wild theories is to axiomatize the behavior of certain formulas using modal logic. Provability logic is perhaps the most famous instance of this phenomenon. First-order $\mathbf{PA}$ is a wild system. Much research in the meta-theory of $\mathbf{PA}$ concerns the behavior of the provability predicate $Prov_{\mathsf{PA}}(\cdot)$ within $\mathbf{PA}$. Solovay showed that this fragment of $\mathbf{PA}$ is tame, in the following sense:
Theorem 6.5 ([solovay1976provability]).
The behavior of the predicate $Prov_{\mathsf{PA}}(\cdot)$ within $\mathbf{PA}$ is axiomatized by the decidable modal logic $\mathbf{GL}$.
The present project is another project in this spirit, although with a focus on definability rather than axiomatization. Kripke’s work of truth is straightforwardly formalizable in second-order arithmetic $\mathbf{PA_{2}}$ or in set theory $\mathbf{ZFC}$, but these theories are firmly on the wild side of the tame/wild divide. By isolating a modal fragment we are able to classify definable sets in a manner analogous to o-minimality (every definable set is a finite union of sets of a very special form). As a consequence we see that—though $\mathbf{S5}$ precludes some interesting phenomena—it permits very sharp results.
[…]

### v3

@@
Abstract
In Outline of a Theory of Truth, Kripke introduces many of the central concepts of the logical study of truth and paradox. He informally defines some of these—such as groundedness and paradoxicality—using modal locutions. We introduce a modal language for regimenting these informal definitions. Though groundedness and paradoxicality are expressible in the modal language, we prove that intrinsicality—which Kripke emphasizes but does not define modally—is not. This follows from a characterization of the modally definable sets and relations and an attendant axiomatization of the modal semantics.
1 Introduction
@@
Yet Kripke does not describe the intrinsic sentences using modal locutions. This is a striking contrast. Is it a mere coincidence that modal locutions appear in the description of the grounded and paradoxical sentences but not of the intrinsic sentences? Might it instead reflect some inherent discrepancy between the sets being defined?
The standard way of studying the relevant sets of sentences obviates this question. In particular, Kripke replaces the informal modal glosses on paradoxicality and groundedness with official extensional mathematical definitions. Nevertheless, the modal glosses are nice examples of modality occurring naturally in a logico-mathematical context, so it is worth regimenting them in a way that retains their modal flavor. For the purposes of taking these modal definitions “at face value,” we introduce a modal language and an attendant interpretation $\mathbf{FIX}$. The modal language is propositional, but it contains structured atomic formulas (like in first-order logic) as opposed to propositional variables. In particular, there are atomic formulas of the form:
@@
Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting.
The aforementioned interpretation $\mathbf{FIX}$ of the modal language interprets $\Box$ via unrestricted universal quantification over a set, [^2] In particular, $\Box$ is interpreted as a universal quantifier over the set of fixed points of the Kripke jump. By ‘unrestricted’ I just mean that the range of the quantifier is not constrained by an accessibility relation. whence it satisfies all principles of the normal modal logic $\mathbf{S5}$. In fact, $\mathbf{FIX}$ satisfies various additional principles (note that $N(x)$ is an abbreviation for $\neg T(x)\wedge\neg F(x)$):
@@
Let $\mathbf{S5[Con,Ground,Min_{n}]}$ be the normal modal logic that results from adjoining these axioms to $\mathbf{S5}$. The following fine-grained soundness and completeness theorem characterizes the modal theory of $\mathbf{FIX}$:
@@
This completeness theorem engenders an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$. From this characterization, we extract the following answer to our motivating question:
Theorem.
The set of intrinsic sentences is not modally definable.
Thus, it is not a mere coincidence that Kripke does not lapse into modal language when defining intrinsicality.
Here is the plan for the rest of the paper. In §2 we review the relevant features of Kripke’s theory of truth. In §3 we introduce the modal language, semantics, and axiom systems that will concern us throughout the paper. In §4 we characterize the sets that are definable using modal formulas with a single variable. In §5 we characterize the relations that are definable using modal formulas with multiple variable.
2 Kripke’s Theory of Truth
@@
2.1 Strong Kleene Logic
@@ 2.2 Classification of Sentences
  • $\lambda$ says $\neg\mathsf{True}(\ulcorner\lambda\urcorner)$.
  • $\tau$ says $\mathsf{True}(\ulcorner\tau\urcorner)$.
  • $\gamma$ says $\mathsf{True}(\ulcorner\gamma\urcorner)\vee\neg\mathsf{True}(\ulcorner\gamma\urcorner)$.
@@ 3 Modal Systems
4 Single Variable Definability
@@ 4.5 Indiscernibility
    $\psi_{1}(x):=\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
@@
    $\psi_{2}(x):=\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
@@
    $\psi_{3}(x):=\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\lozenge N(x)$
@@
    $\psi_{4}(x):=\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\neg\lozenge N(x)$
@@
    $\psi_{5}(x):=\neg\lozenge T(x)\wedge\lozenge F(x)\wedge\neg\lozenge N(x)$
@@
    $\psi_{6}(x):=\neg\lozenge T(x)\wedge\neg\lozenge F(x)\wedge\lozenge N(x)$
Note that the set of inevitable sentences is defined by the disjunction $\psi_{2}(x)\vee\psi_{3}(x)\vee\psi_{4}(x)\vee\psi_{5}(x)$.
Suppose, for the sake of contradiction, that (2) fails. Then there is some $\star$ such that $y^{\star}=\psi$ is inevitable and that $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(y)$. So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., $\psi_{2}(x)$,…,$\psi_{4}(x)$. Let’s break into cases.
Case 1: $\varphi$ is missing disjunct $\psi_{2}$. But $\gamma$ is intrinsic and $\psi_{2}(\gamma)$. So if $x^{\star}$ is $\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 2: $\varphi$ is missing disjunct $\psi_{3}$. But $\neg\gamma$ is intrinsic and $\psi_{3}(\neg\gamma)$. So if $x^{\star}$ is $\neg\gamma$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 3: $\varphi$ is missing disjunct $\psi_{4}$. But $0=0$ is intrinsic and $\psi_{4}(0=0)$. So if $x^{\star}$ is $0=0$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$.
Case 4: $\varphi$ is missing disjunct $\psi_{5}$. But $0=1$ is intrinsic and $\psi_{5}(0=1)$. So if $x^{\star}$ is $0=1$, then $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(x)$, contra $\oplus$. ∎
@@ 5.1 Counting Formulas
5.1.1 The Bare Modal System
@@
5.1.2 With the Consistency Axiom
@@ 5.2 Matrices
Table 1: All extensional 2-isolators.
@@
Table 2: Standard numbering of the $3\times 3$ matrix.
@@
Table 3: A proper subset of the extensional 2-isolators.
@@
Table 4: A proper subset of the extensional 2-isolators.
@@
Table 5: A proper subset of the extensional 2-isolators.
@@
Table 6: A proper subset of the extensional 2-isolators.
@@ 5.5 Completeness
5.6 Counting Definable Relations